# Fourier-RWKV: A Multi-State Perception Network for Efficient Image Dehazing

### v1

@@ Abstract
Image dehazing is crucial for reliable visual perception, yet it remains highly challenging under real-world non-uniform haze conditions. Although Transformer-based methods excel at capturing global context, their quadratic computational complexity hinders real-time deployment. To address this, we propose Fourier Receptance Weighted Key Value (Fourier-RWKV), a novel dehazing framework based on a Multi-State Perception paradigm. The model achieves comprehensive haze degradation modeling with linear complexity by synergistically integrating three distinct perceptual states: (1) Spatial-form Perception, realized through the Deformable Quad-directional Token Shift (DQ-Shift) operation, which dynamically adjusts receptive fields to accommodate local haze variations; (2) Frequency-domain Perception, implemented within the Fourier Mix block, which extends the core WKV attention mechanism of RWKV from the spatial domain to the Fourier domain, preserving the long-range dependencies essential for global haze estimation while mitigating spatial attenuation; (3) Semantic-relation Perception, facilitated by the Semantic Bridge Module (SBM), which utilizes Dynamic Semantic Kernel Fusion (DSK-Fusion) to precisely align encoder-decoder features and suppress artifacts. Extensive experiments on multiple benchmarks demonstrate that Fourier-RWKV delivers state-of-the-art performance across diverse haze scenarios while significantly reducing computational overhead, establishing a favorable trade-off between restoration quality and practical efficiency. Code is available at: https://github.com/Dilizlr/Fourier-RWKV.
Index Terms:
@@
I Introduction
@@
Robust visual perception systems play a critical role in applications such as autonomous driving [10285415, 10753651] and surveillance [10638100, 10916737], with their performance largely dependent on the quality of the input images. Haze under adverse weather conditions severely impairs images, resulting in color distortion, blurred textures, and reduced contrast. Furthermore, spatially non-uniform haze increases the complexity of restoration due to localized variations in density. Therefore, effective dehazing algorithms must balance local adaptability with global consistency to restore image fidelity.
@@
Figure 1: Improvement of our Fourier-RWKV over the SOTA approaches. The bubble size represents the number of model parameters. All models are tested on SOTS-Outdoor [li2018benchmarking].
@@
Early image dehazing methods [5567108, 7128396, 7984895, 8101508] rely on handcrafted priors to estimate transmission maps, but often fail to generalize in real-world settings due to their simplified assumptions. With the rise of deep learning, convolutional neural networks (CNNs) [2016DehazeNet, 10651326, 9919385] have significantly enhanced dehazing performance by automatically learning haze-related features from large-scale data. However, the limited receptive fields of CNNs hinder the modeling of long-range dependencies, which are crucial for accurate global haze estimation, especially in complex scenes. By leveraging the self-attention mechanism, Transformers [10.1145/3664647.3681314, 10076399, Yang2023, wang2024gridformer] exhibit superior capability in capturing global context, driving remarkable progress in image dehazing. Nevertheless, their quadratic computational complexity incurs high computational costs on high-resolution images, restricting their real-time applicability.
@@
To overcome the efficiency bottleneck, researchers have begun exploring attention architectures with linear complexity, such as State Space Models (SSMs) [DBLPconficlrGuGR22, DBLPconfnipsGuJGSDRR21, DBLPjournalscorrabs-2312-00752] and the RWKV model [DBLPconfemnlpPengAAAABCCCDDG23, DBLPjournalscorrabs-2404-05892]. Among them, Vision-RWKV has demonstrated promise for efficient long-range modeling in vision tasks. However, its direct application to image dehazing encounters several key architectural limitations: (1) The spatial operations in Vision-RWKV [DBLPconficlrDuanWCZLLQ0DW25], such as the fixed Q-Shift, lack the adaptability required to capture irregular haze distributions; (2) The sequential state modeling in the spatial domain, driven by the WKV attention mechanism, suffers from long-range information decay, undermining both accuracy and consistency in global haze estimation; (3) When integrated into the U-Net encoder–decoder framework, the inherent semantic gap between encoding and decoding stages causes feature misalignment and artifacts—an issue that remains unresolved in the original Vision-RWKV.
@@
First, we introduce the deformable quad-directional token shift (DQ-Shift) operation, which performs dynamic receptive field adjustment to adapt to local haze variations, overcoming the rigidity of fixed spatial operations. Second, we present the Fourier Mix block, which transforms the WKV attention mechanism to the Fourier domain to facilitate efficient long-range modeling. This design is motivated by two key insights (see Fig. 2): (1) The Fourier representation allows for effective separation of haze interference (primarily in amplitude) from structural content (preserved in phase); (2) The global statistical properties of the Fourier domain naturally establish long-range dependencies, mitigating the information decay in spatial-domain modeling. Furthermore, we propose the Semantic Bridge Module (SBM) to address the inherent semantic gap within the encoder-decoder structure. SBM employs the Dynamic Semantic Kernel Fusion (DSK-Fusion) mechanism to align features between corresponding stages, ensuring semantic consistency throughout the dehazing process.
@@
II Related Work
II-A Prior-based and CNN-based image dehazing
@@
Single-image dehazing is a highly ill-posed problem. Early works simplify this challenge by introducing handcrafted priors [5567108, 7128396, 7984895, 8101508, Berman_2016_CVPR, 2651362] to constrain the solution space. While effective in simpler scenarios, their rigid assumptions struggle to capture the complexity and variability of real-world haze, leading to poor performance in challenging environments.
@@
The introduction of CNNs marks a paradigm shift from fixed priors to data-driven feature learning, substantially enhancing adaptability across diverse conditions. Early models, such as DehazeNet [2016DehazeNet], AODNet [8237773], and DCPDN [8578435], estimate transmission maps and atmospheric light based on the physical scattering model [1977Optics, 2002Vision, 2016Haze]. Later methods, including GridDehazeNet [Liu_2019_ICCV], FFANet [qin2020ffa], and DEA-Net [10411857], adopt end-to-end learning strategies to directly restore haze-free images. This transition reduces reliance on physical priors, improving output fidelity. However, the limited receptive fields of convolutions hinder global context modeling, making it challenging to balance global haze estimation and fine detail recovery under non-uniform haze distributions.
II-B Transformer-based image dehazing
@@
Transformer-based models [9879191, wu2024adaptive, 10678109, wen2025all] have advanced the state of the art in image dehazing by effectively modeling long-range dependencies through the self-attention mechanism, greatly improving restoration quality. However, their quadratic computational complexity relative to input resolution limits their applicability in high-resolution and real-time scenarios, prompting a surge of efficiency-oriented designs.
@@
Swin Transformer [DBLPconficcvLiuL00W0LG21] reduces computational cost by limiting self-attention to non-overlapping local windows. DehazeFormer [10076399] and Trinity-Net [10149032] extend this approach to image dehazing, achieving competitive performance with lower overhead. SwinIR [DBLPconficcvwLiangCSZGT21] further introduces a shifting mechanism for cross-window interactions, mitigating artifact generation. Alternatively, MB-TaylorFormer [10378631, 10962317] approximates self-attention via Taylor expansion, and Restormer [Zamir_2022_CVPR] replaces spatial attention with channel attention, both strategies effectively alleviating the computational burden. Despite these developments, most of these methods remain confined to the spatial domain or rely on approximations of self-attention, inevitably sacrificing global modeling accuracy for efficiency. This trade-off has motivated the exploration of more fundamental architectural innovations.
II-C Frequency domain learning
@@
Frequency-domain learning transforms images into the frequency domain to exploit the global statistics of the spectrum, enabling global dependency modeling with linear complexity and overcoming the efficiency bottleneck inherent in spatial-domain approaches. Moreover, the decoupled and structured nature of frequency-domain representations simplifies the learning of semantically coherent structures from highly entangled local pixels, as different frequency bands naturally correspond to distinct semantic components (e.g., low frequencies for contours and high frequencies for textures). These advantages have rendered frequency-domain learning attractive for fundamental vision tasks, including image deblurring [10204001], low-light enhancement [DBLPconficlrLiGZLZFL23], and super-resolution [DBLPconfmmFangZWZCH22].
@@
The spectral separability of haze degradation (see Fig. 2) further offers a powerful physical prior for dehazing. Leveraging this property, recent frequency-domain dehazing methods [10678201, CUI2025111074, DBLPjournalscorrabs-2507-11035, gao2023frequency] perform targeted manipulation of frequency components to effectively disentangle haze from underlying image structures. These methods deliver improved dehazing performance with markedly reduced overhead.
@@
II-D Receptance Weighted Key Value
The pursuit of efficient architectures has also driven the rise of linear attention models as viable alternatives to Transformers. Among these, the Receptance Weighted Key Value (RWKV) model [DBLPconfemnlpPengAAAABCCCDDG23, DBLPjournalscorrabs-2404-05892], originally designed for NLP, offers notable advantages. Its core WKV mechanism facilitates long-range modeling with linear complexity, while token-shifting enhances local perception. Vision-RWKV [DBLPconficlrDuanWCZLLQ0DW25] extends this framework to vision tasks by reformulating WKV attention for two-dimensional, non-causal data. It introduces a bidirectional variant (Bi-WKV) to capture global dependencies, while the quad-directional token shift (Q-Shift) aggregates multi-directional local context. Several variants have been developed for specific vision tasks, including URWKV [DBLPconfcvpr0028NLXLC25] for low-light enhancement, CRWKV [DBLPconfijcaiChenCJXZW25] for image denoising, Restore-RWKV [DBLPjournalscorrabs-2407-11087] for medical image restoration, and BSBP-RWKV [DBLPconfmmZhouC24] for medical image segmentation.
III Method
III-A Overall Architecture
As illustrated in Fig. 3, Fourier-RWKV employs a symmetric encoder-decoder architecture. Given an input hazy image $I\in\mathbb{R}^{3\times H\times W}$, the network initially extracts shallow features using a $3\times 3$ convolution. These features are processed through an encoder-decoder structure with four resolution levels, each containing multiple FRWKV blocks for feature transformation and reconstruction across scales. The output feature $I_{recon}$ is further refined via another $3\times 3$ convolution to produce a residual image $I_{r}\in\mathbb{R}^{3\times H\times W}$. The final dehazed result is then obtained as $I_{c}=I+I_{r}$.
@@ III-A Overall Architecture
III-B FRWKV Block: The Building Block
@@
III-B1 Deformable Quad-directional Token Shift
@@
III-B2 Fourier Mix Block with DQ-Shift
@@
where $\mathrm{CIR}$ refers to channel-wise concatenation of real and imaginary parts, yielding $K_{fft},V_{fft}\in\mathbb{R}^{2C\times H\times\lceil\frac{W}{2}\rceil}$. Here, $\lceil\cdot\rceil$ denotes the round-up operator.
@@
where $\mathrm{Seq}$ refers to the operation that flattens the features into a sequence ordered by $M$, resulting in $K_{fft},V_{fft}\in\mathbb{R}^{2C\times T}$ with sequence length $T=H\times\lceil\frac{W}{2}\rceil$.
@@
where wkvwkv denotes the attention weights from Bi-WKV, $\mathrm{ISeq}$ restores the original frequency grid structure using the index grid $M$, $\mathrm{ICIR}$ reverses channel-wise concatenation to reconstruct complex frequency representations, and $\mathcal{F}^{-1}$ represents the inverse FFT. The final output $O_{s}$ is obtained by fusing the spatial gate with the Fourier-modulated feature.
III-B3 Channel Mix Block with DQ-Shift
@@
III-C Semantic Bridge Module
@@
To tackle this issue, we propose the semantic bridge module (SBM), a lightweight fusion module that semantically aligns corresponding encoder and decoder features. The SBM structure is illustrated in Fig. 4. At its core lies the dynamic semantic kernel fusion (DSK-Fusion) mechanism, which generates channel-adaptive convolution kernels by explicitly modeling inter-feature semantic correlations.
@@
where each element $\mathbf{Sim}_{i,j}$ indicates the semantic correlation between the $i$-th channel of the encoder and the $j$-th channel of the decoder. Given that haze degradation is concentrated in low-frequency components (Fig. 2), the similarity matrix also reflects inter-channel correlations in degradation patterns.
@@
TABLE I: Quantitative comparisons with state-of-the-art dehazing methods on the synthetic and real-world datasets.
@@
GridDehazeNet [Liu_2019_ICCV] ICCV’19 32.16 0.984 30.86 0.982 13.31 0.37 13.80 0.54 0.956 21.49
MSBDN [dong2020multi] CVPR’20 33.67 0.985 33.48 0.982 15.37 0.49 19.23 0.71 31.35 41.54
FFA-Net [qin2020ffa] AAAI’20 36.39 0.989 33.57 0.984 14.39 0.45 19.87 0.69 4.456 287.8
PMNet [ye2022perceiving] ECCV’22 38.41 0.990 34.74 0.985 16.79 0.51 20.42 0.73 18.90 81.13
MAXIM-2S [tu2022maxim] ICCV’22 38.11 0.991 34.19 0.985 - - - - 14.10 216.0
@@
Fourmer [zhou2023fourmer] ICML’23 37.32 0.990 - - 15.95 0.49 - - 1.29 20.6
@@
OKNet [cui2024omni] AAAI’24 40.79 0.996 37.68 0.995 16.92 0.64 20.48 0.80 4.72 39.71
CSNet [cui2024hybrid] IJCAI’24 41.34 0.996 - - 17.33 0.65 20.43 0.80 4.27 41.19
MAIR [DBLPconfmmFangZWZCH22] CVPR’25 39.45 0.997 36.96 0.991 - - - - 3.40 24.03
PGH2Net [su2025prior] AAAI’25 41.70 0.996 37.52 0.989 17.02 0.61 - - 1.76 16.05
@@
III-D Loss Function
IV Experiments
IV-A Experimental settings
Datasets. To thoroughly evaluate the dehazing performance and generalization ability of the proposed Fourier-RWKV, we conduct experiments on both widely-used synthetic datasets and challenging real-world datasets. For synthetic data, we utilize the RESIDE [li2018benchmarking] benchmark dataset, which covers indoor and outdoor scenes with various haze conditions. Specifically, the model is independently trained on the indoor training subset (ITS, containing 13,990 image pairs) and the outdoor training subset (OTS, containing 313,950 image pairs), with performance evaluated on 500 indoor and 500 outdoor scenes from the SOTS test set, respectively. For real-world data, we select two representative haze datasets, Dense-Haze [ancuti2019dense] and NH-HAZE [9150807]. Dense-Haze focuses on dense haze conditions, while NH-HAZE deals with non-uniform haze distributions. Both datasets consist of 55 pairs of hazy and clear images. The first 50 pairs from each dataset are used for training, and the remaining 5 pairs are reserved for testing, ensuring consistent and comparable results.

### v2

@@ Abstract
Image dehazing is crucial for reliable visual perception, yet it remains highly challenging under real-world non-uniform haze conditions. Although Transformers excel at capturing global context, their quadratic computational complexity hinders real-time deployment. To address this, we propose Fourier Receptance Weighted Key Value (Fourier-RWKV), a novel dehazing framework based on a Multi-State Perception paradigm. The model achieves comprehensive haze degradation modeling with linear complexity by synergistically integrating three distinct perceptual states: (1) Spatial-form Perception, realized through the Deformable Quad-directional Token Shift (DQ-Shift) operation, which dynamically adjusts receptive fields to accommodate local haze variations; (2) Frequency-domain Perception, implemented within the Fourier Mix block, which extends the core WKV attention mechanism of RWKV from the spatial domain to the Fourier domain, preserving the long-range dependencies essential for global haze estimation while mitigating spatial attenuation; (3) Semantic-relation Perception, facilitated by the Semantic Bridge Module (SBM), which utilizes Dynamic Semantic Kernel Fusion (DSK-Fusion) to precisely align encoder-decoder features and suppress artifacts. Extensive experiments on multiple benchmarks demonstrate that Fourier-RWKV delivers state-of-the-art performance across diverse haze scenarios while significantly reducing computational overhead, establishing a favorable trade-off between restoration quality and practical efficiency. Code is available at: https://github.com/Dilizlr/Fourier-RWKV.
keywords:
@@
journal: Nuclear Physics B
[…]
[1]organization=Institute of Intelligent Information Processing, Guangdong Provincial Key Laboratory of Intelligent Information Processing, Shenzhen Key Laboratory of Modern Communications and Information Processing, Shenzhen University, city=Shenzhen, postcode=518060, country=China
[…]
[2]organization=College of Engineering and Computer Science, Australian National University, city=Canberra, postcode=2601, state=ACT, country=Australia
1 Introduction
@@
Robust visual perception systems play a critical role in applications such as autonomous driving [37, 24] and person re-identification [59, 65], with their performance largely dependent on the quality of the input images. Haze under adverse weather conditions severely impairs images, resulting in color distortion, blurred textures, and reduced contrast. Furthermore, spatially non-uniform haze increases the complexity of restoration due to localized variations in density. Therefore, effective dehazing algorithms must balance local adaptability with global consistency to restore image fidelity.
@@
Figure 1: Improvement of our Fourier-RWKV over the SOTA approaches. The bubble size represents the number of model parameters. All models are tested on (a) SOTS-Indoor and (b) SOTS-Outdoor datasets.
@@
Early image dehazing methods [25, 64, 48, 5] rely on handcrafted priors to estimate transmission maps, but often fail to generalize in real-world settings due to their simplified assumptions. With the rise of deep learning, convolutional neural networks (CNNs) [6, 61, 49] have significantly enhanced dehazing performance by automatically learning haze-related features from large-scale data. However, the limited receptive fields of CNNs hinder the modeling of long-range dependencies, which are crucial for accurate global haze estimation, especially in complex scenes. By leveraging the self-attention mechanism, Transformers [35, 45, 55, 50] exhibit superior capability in capturing global context, driving remarkable progress in image dehazing. Nevertheless, their quadratic computational complexity incurs high computational costs on high-resolution images, restricting their real-time applicability.
@@
To overcome the efficiency bottleneck, researchers have begun exploring attention architectures with linear complexity, such as State Space Models (SSMs) [21, 22, 20] and the RWKV model [41, 42]. Among them, Vision-RWKV has demonstrated promise for efficient long-range modeling in vision tasks. However, its direct application to image dehazing encounters several key architectural limitations: (1) The spatial operations in Vision-RWKV [16], such as the fixed Q-Shift, lack the adaptability required to capture irregular haze distributions; (2) The sequential state modeling in the spatial domain, driven by the WKV attention mechanism, suffers from long-range information decay, undermining both accuracy and consistency in global haze estimation; (3) When integrated into the U-Net encoder–decoder framework, the inherent semantic gap between encoding and decoding stages causes feature misalignment and artifacts—an issue that remains unresolved in the original Vision-RWKV.
@@
First, we introduce the deformable quad-directional token shift (DQ-Shift) operation, which performs dynamic receptive field adjustment to adapt to local haze variations, overcoming the rigidity of fixed spatial operations. Second, we present the Fourier Mix block, which transforms the WKV attention mechanism to the Fourier domain to facilitate efficient long-range modeling. This design is motivated by two key insights (see Figure 2): (1) The Fourier representation allows for effective separation of haze interference (primarily in amplitude) from structural content (preserved in phase); (2) The global statistical properties of the Fourier domain naturally establish long-range dependencies, mitigating the information decay in spatial-domain modeling. Furthermore, we propose the Semantic Bridge Module (SBM) to address the inherent semantic gap within the encoder-decoder structure. SBM employs the Dynamic Semantic Kernel Fusion (DSK-Fusion) mechanism to align features between corresponding stages, ensuring semantic consistency throughout the dehazing process.
@@
2 Related Work
2.1 Prior-based and CNN-based image dehazing
@@
Single-image dehazing is a highly ill-posed problem. Early works simplify this challenge by introducing handcrafted priors [25, 64, 48, 5, 4, 18] to constrain the solution space. While effective in simpler scenarios, their rigid assumptions struggle to capture the complexity and variability of real-world haze, leading to poor performance in challenging environments.
@@
The introduction of CNNs marks a paradigm shift from fixed priors to data-driven feature learning, substantially enhancing adaptability across diverse conditions. Early models, such as DehazeNet [6], AODNet [28], and DCPDN [58], estimate transmission maps and atmospheric light based on the physical scattering model [38, 39, 32]. Later methods, including GridDehazeNet [34], FFANet [43], and DEA-Net [8], adopt end-to-end learning strategies to directly restore haze-free images. This transition reduces reliance on physical priors, improving output fidelity. However, the limited receptive fields of convolutions hinder global context modeling, making it challenging to balance global haze estimation and fine detail recovery under non-uniform haze distributions.
2.2 Transformer-based image dehazing
@@
Transformer-based models [23, 53, 15, 52] have advanced the state of the art in image dehazing by effectively modeling long-range dependencies through the self-attention mechanism, greatly improving restoration quality. However, their quadratic computational complexity relative to input resolution limits their applicability in high-resolution and real-time scenarios, prompting a surge of efficiency-oriented designs.
@@
Swin Transformer [36] reduces computational cost by limiting self-attention to non-overlapping local windows. DehazeFormer [45] and Trinity-Net [9] extend this approach to image dehazing, achieving competitive performance with lower overhead. SwinIR [33] further introduces a shifting mechanism for cross-window interactions, mitigating artifact generation. Alternatively, MB-TaylorFormer [44, 26] approximates self-attention via Taylor expansion, and Restormer [57] replaces spatial attention with channel attention, both strategies effectively alleviating the computational burden. Despite these developments, most of these methods remain confined to the spatial domain or rely on approximations of self-attention, inevitably sacrificing global modeling accuracy for efficiency. This trade-off has motivated the exploration of more fundamental architectural innovations.
2.3 Frequency domain learning
@@
Frequency-domain learning transforms images into the frequency domain to exploit the global statistics of the spectrum, enabling global dependency modeling with linear complexity and overcoming the efficiency bottleneck inherent in spatial-domain approaches. Moreover, the decoupled and structured nature of frequency-domain representations simplifies the learning of semantically coherent structures from highly entangled local pixels, as different frequency bands naturally correspond to distinct semantic components (e.g., low frequencies for contours and high frequencies for textures). These advantages have rendered frequency-domain learning attractive for fundamental vision tasks, including image deblurring [27], low-light enhancement [31], and super-resolution [17].
@@
The spectral separability of haze degradation (see Figure 2) further offers a powerful physical prior for dehazing. Leveraging this property, recent frequency-domain dehazing methods [40, 13, 60, 19] perform targeted manipulation of frequency components to effectively disentangle haze from underlying image structures. These methods deliver improved dehazing performance with markedly reduced overhead.
2.4 Receptance Weighted Key Value
The pursuit of efficient architectures has also driven the rise of linear attention models as viable alternatives to Transformers. Among these, the Receptance Weighted Key Value (RWKV) model [41, 42], originally designed for NLP, offers notable advantages. Its core WKV mechanism facilitates long-range modeling with linear complexity, while token-shifting enhances local perception. Vision-RWKV [16] extends this framework to vision tasks by reformulating WKV attention for two-dimensional, non-causal data. It introduces a bidirectional variant (Bi-WKV) to capture global dependencies, while the quad-directional token shift (Q-Shift) aggregates multi-directional local context. Several variants have been developed for specific vision tasks, including URWKV [54] for low-light enhancement, CRWKV [7] for image denoising, Restore-RWKV [3] for medical image restoration, and BSBP-RWKV [63] for medical image segmentation.
3 Method
3.1 Overall Architecture
As illustrated in Figure 3, Fourier-RWKV employs a symmetric encoder-decoder architecture. Given an input hazy image $I\in\mathbb{R}^{3\times H\times W}$, the network initially extracts shallow features using a $3\times 3$ convolution. These features are processed through an encoder-decoder structure with four resolution levels, each containing multiple FRWKV blocks for feature transformation and reconstruction across scales. The output feature $I_{recon}$ is further refined via another $3\times 3$ convolution to produce a residual image $I_{r}\in\mathbb{R}^{3\times H\times W}$. The final dehazed result is then obtained as $I_{c}=I+I_{r}$.
@@ 3.1 Overall Architecture
3.2 FRWKV Block: The Building Block
@@
3.2.1 Deformable Quad-directional Token Shift
@@
3.2.2 Fourier Mix Block with DQ-Shift
@@
where $\mathrm{CIR}(\cdot)$ refers to channel-wise concatenation of real and imaginary parts, yielding $K_{fft},V_{fft}\in\mathbb{R}^{2C\times H\times\lceil\frac{W}{2}\rceil}$. Here, $\lceil\cdot\rceil$ denotes the round-up operator.
@@
where $\mathrm{Seq}(\cdot)$ refers to the operation that flattens the features into a sequence ordered by $M$, resulting in $K_{fft},V_{fft}\in\mathbb{R}^{2C\times T}$ with sequence length $T=H\times\lceil\frac{W}{2}\rceil$.
@@
where wkvwkv denotes the attention weights from Bi-WKV, $\mathrm{ISeq}(\cdot)$ restores the original frequency grid structure using the index grid $M$, $\mathrm{ICIR}(\cdot)$ reverses channel-wise concatenation to reconstruct complex frequency representations, and $\mathcal{F}^{-1}$ represents the inverse FFT. The final output $O_{s}$ is obtained by fusing the spatial gate with the Fourier-modulated feature.
3.2.3 Channel Mix Block with DQ-Shift
@@
3.3 Semantic Bridge Module
@@
To tackle this issue, we propose the semantic bridge module (SBM), a lightweight fusion module that semantically aligns corresponding encoder and decoder features. The SBM structure is illustrated in Figure 4. At its core lies the dynamic semantic kernel fusion (DSK-Fusion) mechanism, which generates channel-adaptive convolution kernels by explicitly modeling inter-feature semantic correlations.
@@
where each element $\mathbf{Sim}_{i,j}$ indicates the semantic correlation between the $i$-th channel of the encoder and the $j$-th channel of the decoder. Given that haze degradation is concentrated in low-frequency components (Figure 2), the similarity matrix also reflects inter-channel correlations in degradation patterns.
@@
3.4 Loss Function
Table 1: Quantitative comparisons with state-of-the-art dehazing methods on the synthetic and real-world datasets.
@@ 3.4 Loss Function
GridDehazeNet [34] ICCV’19 32.16 0.984 30.86 0.982 13.31 0.37 13.80 0.54 0.956 21.49
MSBDN [14] CVPR’20 33.67 0.985 33.48 0.982 15.37 0.49 19.23 0.71 31.35 41.54
FFA-Net [43] AAAI’20 36.39 0.989 33.57 0.984 14.39 0.45 19.87 0.69 4.456 287.8
PMNet [56] ECCV’22 38.41 0.990 34.74 0.985 16.79 0.51 20.42 0.73 18.90 81.13
MAXIM-2S [47] ICCV’22 38.11 0.991 34.19 0.985 - - - - 14.10 216.0
@@
Fourmer [62] ICML’23 37.32 0.990 - - 15.95 0.49 - - 1.29 20.6
@@
OKNet [12] AAAI’24 40.79 0.996 37.68 0.995 16.92 0.64 20.48 0.80 4.72 39.71
CSNet [10] IJCAI’24 41.34 0.996 - - 17.33 0.65 20.43 0.80 4.27 41.19
MAIR [30] CVPR’25 39.45 0.997 36.96 0.991 - - - - 3.40 24.03
PGH2Net [46] AAAI’25 41.70 0.996 37.52 0.989 17.02 0.61 - - 1.76 16.05
@@
4 Experiments
4.1 Experimental settings
Datasets. To thoroughly evaluate the dehazing performance and generalization ability of the proposed Fourier-RWKV, we conduct experiments on both widely-used synthetic datasets and challenging real-world datasets. For synthetic data, we utilize the RESIDE [29] benchmark dataset, which covers indoor and outdoor scenes with various haze conditions. Specifically, the model is independently trained on the indoor training subset (ITS, containing 13,990 image pairs) and the outdoor training subset (OTS, containing 313,950 image pairs), with performance evaluated on 500 indoor and 500 outdoor scenes from the SOTS test set, respectively. For real-world data, we select two representative haze datasets, Dense-Haze [1] and NH-HAZE [2]. Dense-Haze focuses on dense haze conditions, while NH-HAZE deals with non-uniform haze distributions. Both datasets consist of 55 pairs of hazy and clear images. The first 50 pairs from each dataset are used for training, and the remaining 5 pairs are reserved for testing, ensuring consistent and comparable results.